Income Or Output Y Consumption Expenditure         C Investment Expenditure       I Government Expenditure          G...

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

$4,000

3,925

100

100

25

4,100

4,000

100

100

25

4,200

4,075

100

100

25

4,300

4,150

100

100

25

4,400

4,225

100

100

25

4,500

4,300

100

100

25

4,600

4,375

100

100

25

4,700

4,450

100

100

25

4,800

4,525

100

100

25

4,900

4,600

100

100

25

5,000

4,675

100

100

25

a) Determine equilibrium level of income?
Answer #1

Equilibrium level of income is where

Y = C+ I+ G+NX

And according to the given table

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

C +I + G + NX

$4,000

3,925

100

100

25

3925 + 100 + 100 + 25 = 4150

4,100

4,000

100

100

25

4000 + 100 + 100+ 25 = 4225

4,200

4,075

100

100

25

4300

4,300

4,150

100

100

25

4375

4,400

4,225

100

100

25

4450

4,500

4,300

100

100

25

4525

4,600

4,375

100

100

25

4600

4,700

4,450

100

100

25

4675

4,800

4,525

100

100

25

4750

4,900

4,600

100

100

25

4825

5,000

4,675

100

100

25

4900

At income level 4600,

Y =C + I + G + NX

So equilibrium level off income is $4600